What is the boiling point of a 0.66 molal solution of sucrose (nonelectrolyte) in ethanol? (Kb of ethanol = 1.22 °C/m, normal boiling point of ethanol is 78.5 °C)
1
Expert's answer
2020-06-04T10:16:24-0400

Decrease of boiling point after sucrose dissolution is ΔT = K*m = 1.22*0.66 = 0.8 °C

Boiling point of solution is T = 78.5 - 0.8 = 77.7 °C